# Bose–Fermi $N$-polaron state emergence from correlation-mediated blocking of phase separation

### v2

@@ Abstract
We first explore the case of Open(B)-Open(F) BCs where phase separation occurs, for which we show the von Neumann entropy and mutual information in Fig. 3(a) and (b) as a function of $U_{\rm BF}$. Starting from the decoupled pure states at $U_{\rm BF}=0$, $S_{\rm BF}$ increases with increasing $U_{\rm BF}$ as bosonic and fermionic states mix, reaching its maximum around the critical interaction strength $U_{\rm BF}\approx U_{\rm BB}$. In Fig. 1(a-c) we indicate the coupling strengths that give the maximum entropy with green lines, showing that it exactly separates the two phases and only shows very small deviations from $U_{\rm BF}=U_{\rm BB}$ below $U_{\rm BB,0}^{c}$. On the contrary, the mutual information $I_{\rm FF}$ decreases till the miscible-immiscible transition occurs, indicating impurity-impurity correlation screening as the impurities move to the system’s edges because of correlations with the bath, reaching a minimum near the $S_{\rm BF}$ maximum  [^2] For small $U_{\rm BB}$, the $S_{\rm BF}$ and $I_{\rm FF}$ extrema slightly differ but coincide as $U_{\rm BB}$ increases.. Increasing $U_{\rm BF}$ beyond this critical point leads to a decay in $S_{\rm BF}$ while $I_{\rm FF}$ returns to its initial value. These are both indicators of the phase separation regime in which the only contribution to the impurity-impurity correlations is the Pauli exclusion principle, while the bosonic bath acts as a double well potential that traps two impurities at the edges. The residual correlation in $S_{\rm BF}$ at high $U_{\rm BF}$ values is larger when the system is deeper in the SF regime ($U_{\rm BB}<U_{\rm BB,0}^{c}$), which can be understood as the compressible nature of the bosonic SF allowing a finite overlap with the fermions even at large $U_{\rm BF}$.
@@
For Closed(B)-Open(F) BCs with weak interspecies coupling, $U_{\rm BF}<U_{\rm BB}$, $S_{\rm BF}$ and $I_{\rm FF}$ exhibit similar behavior as for Open(B)-Open(F) BCs, increasing and decreasing respectively until reaching a critical point that signals the miscible-polaronic transition. Again we take the coupling strength at the maximum of the von Neumann entropy and highlight it by the green lines in Fig. 1(d) and (e), which clearly signals the emergence of the bipolaron state. Moreover, in the inset of Fig. 3(d), we show in a closeup of the mutual information for $U_{\rm BB}=6$ that while in the MI phase (filled by grey background) $I_{\rm FF}$ is unchanged, indicating a lack of impurity-impurity correlations due to the “slow” spread of information through the insulating background [76], until the bath transitions to the SF state (white background). This transition almost coincides with the critical coupling strength obtained by the maximum relative distance between the impurities indicated by the blue vertical line, which is highlighted previously in Fig. 1(d-f). A similar behavior can also be seen in Fig. 3(b) for Open(B)-Open(F) BCs, although in this case, no clear indication of a transition appears in the fermionic density and distance like in the Closed(B)-Open(F) BCs case.
@@
As expected from the differences in density profiles (Fig. 1(a,d)), the behavior changes radically after the miscible-polaron transition occurs where we observe a notable increase in $I_{\rm FF}$ and $S_{\rm BF}$ until both curves converge to large values in the $U_{\rm BF}\to\infty$ limit. Indeed, this indicates that strong impurity-bath correlations are responsible for the increased impurity-impurity correlations, as the former is necessary to induce large attractive interactions between the impurities and thus create the tightly bound states shown in Fig. 2(c). The saturation of the von Neumann entropy locks the correlations between the impurities and the bath, and therefore halts further localization of the fermions as shown in Fig. 1(d). This is an example of impurity self-localization suppression enhanced by impurity-bath correlations, an effect which was proposed in Ref. [78] using a variational ansatz in which the impurity-bath correlations are described nonperturbatively. Our work provides further evidence for this effect in discrete lattice systems, and here it is also responsible for the system to completely avoid the phase-separated state.
Finally, we note that the formation of $N$-polaron states through strong system-bath correlations is only present when there is a high interspecies population imbalance. To show this we analyze the regime of $N_{\rm F}>2$ where, practically speaking, obtaining $S_{\rm BF}$ is only possible for very few impurities since the dimension of $\hat{\rho}_{\rm F}$ increases as $L^{N_{\rm F}}\times L^{N_{\rm F}}$. In this case the bipartite mutual information $I_{\rm FF}$ represents the ideal tool as it signals the presence of bath-induced correlations and the formation of the $N$-polaron state when $I_{\rm FF}>I_{\rm FF,0}$, with $I_{\rm FF,0}$ being the mutual information of two fermions decoupled from the bosonic background ($U_{\rm BF}=0$). We show $I_{\rm FF}$ in Fig. 3(e) as a function of $N_{\rm F}$ and taking the limit $U_{\rm BF}\to\infty$ by removing the double Bose–Fermi occupancy in the Hilbert space for the DMRG calculations [67]. Adding impurities to the system for a fixed $U_{\rm BB}$ value effectively compresses the bath, as the bosons cannot occupy the same place as the fermionic cloud. This inhibits the bath-mediated interactions which reduces $I_{\rm FF}$ until a critical impurity density is reached and $I_{\rm FF}=I_{\rm{FF},0}$ indicating that the ground state is now phase separated. In Fig. 3(f), we show $I_{\rm FF}$ for the phase-separated state at fermionic half-filling $N_{\rm F}=15$, while the inset shows the density distributions at $U_{\rm BB}=6$ and $U_{\rm BF}=24$. In the phase-separated regime, the system exhibits a similar behavior as the Open(B)-Open(F) BCs case (Fig. 3(b)), however, the transition still occurs at larger values of the boson-fermion coupling $(U_{\rm BF}>U_{\rm BB})$. Finally, we note the crucial role that the bosonic interactions $U_{\rm BB}$ have on the critical impurity number. As can be seen from Fig. 3(e), for larger intraspecies interactions, phase separation occurs even for a small number of impurities as boson-boson correlations prevent overlap with the fermionic cloud, analogous to the behavior of $S_{\rm BF}$ at large $U_{\rm BF}$ in Fig. 3(a), therefore hastening the formation of the phase separated state.
@@
We have studied the formation of a $N$-polaron state, where fermionic impurities are tightly bound together by a unit-filled bosonic gas in a one-dimensional optical lattice. We have shown that a superfluid-insulator transition is crucial for polaron formation, as when the Mott phase is broken and the background becomes a superfluid, correlations with the bosonic bath bind fermions together through strong mediated attractive interactions, allowing redistribution of the polaron across the superfluid bath and preventing phase separation, similar to Ref. [78] and recently observed in DMRG simulations [79]. This also suggests that the presence of strong impurity-bath correlations could inhibit the formation of the self-pinned state [80], as the mediated attractive interactions may prevent the impurities from self-organizing into a periodic structure. We have also shown that population imbalance is key for $N$-polaron formation, and that suitable tuning of bath intraspecies interactions and correlations could allow the creation of large and stable strongly-entangled $N$-polaron states. Indeed, a recent similar work has explored how the filling fraction and boson-boson interactions can affect bipolaron formation elucidating complementary results to our own [81]. Future studies could also investigate the role of particle statistics inside the polaronic state, analog to previous work with bosonic impurities using exact diagonalization [48, 51]. Moreover, the use of mixed boundary conditions on other well-known impurity systems could lead to novel phenomena by blocking or enhancing boundary effects.

### v3

@@ V Strong coupling regime
We first explore the case of Open(B)-Open(F) BCs where phase separation occurs, for which we show the von Neumann entropy and mutual information in Fig. 5(a) and (b) as a function of $U_{\rm BF}$. Starting from the decoupled pure states at $U_{\rm BF}=0$, $S_{\rm BF}$ increases with increasing $U_{\rm BF}$ as bosonic and fermionic states mix, reaching its maximum around the critical interaction strength $U_{\rm BF}\approx U_{\rm BB}$. In Fig. 1(a-c) we indicate the coupling strengths that give the maximum entropy with green lines, showing that it exactly separates the two phases and only shows very small deviations from $U_{\rm BF}=U_{\rm BB}$ below $U_{\rm BB,0}^{c}$. On the contrary, the mutual information $I_{\rm FF}$ decreases until the miscible-immiscible transition occurs, indicating impurity-impurity correlation screening due to the increased correlations with the bath, reaching a minimum near the $S_{\rm BF}$ maximum  [^2] For small $U_{\rm BB}$, the $S_{\rm BF}$ and $I_{\rm FF}$ extrema slightly differ but coincide as $U_{\rm BB}$ increases.. Increasing $U_{\rm BF}$ beyond this critical point leads to the fermions moving the systems edges and a decay in $S_{\rm BF}$ while $I_{\rm FF}$ returns to its initial value. These are both indicators of the phase separation regime in which the only contribution to the impurity-impurity correlations is the Pauli exclusion principle, while the bosonic bath acts as an effective double well potential that confines the two impurities to the trap edges. The residual correlation in $S_{\rm BF}$ at high $U_{\rm BF}$ values is larger when the system is deeper in the SF regime ($U_{\rm BB}<U_{\rm BB,0}^{c}$), which can be understood as the compressible nature of the bosonic SF allowing a finite overlap with the fermions even at large $U_{\rm BF}$.
@@
For Periodic(B)-Open(F) BCs (Fig. 6) with weak interspecies coupling, $U_{\rm BF}<U_{\rm BB}$, $S_{\rm BF}$ and $I_{\rm FF}$ exhibit similar behavior as for Open(B)-Open(F) BCs, increasing and decreasing respectively until reaching a critical point that signals the miscible-polaronic transition. Again we take the coupling strength at the maximum of the von Neumann entropy and highlight it by the green lines in Fig. 1(d) and (e), which clearly signals the emergence of the bipolaron state. Moreover, in the inset of Fig. 6(b), we show in a closeup of the mutual information for $U_{\rm BB}=6$ that while in the MI phase (filled by grey background) $I_{\rm FF}$ is unchanged, indicating a lack of impurity-impurity correlations due to the “slow” spread of information through the insulating background [85], until the bath transitions to the SF state (white background). This transition almost coincides with the critical coupling strength obtained by the maximum relative distance between the impurities indicated by the blue vertical line, which is highlighted previously in Fig. 1(d-f).
@@
As expected from the differences in density profiles (Fig. 1(a,d)), the behavior changes radically after the miscible-polaron transition occurs where we observe a notable increase in $I_{\rm FF}$ and $S_{\rm BF}$ until both quantities saturate to large values in the $U_{\rm BF}\to\infty$ limit. Indeed, this indicates that strong impurity-bath correlations are responsible for the increased impurity-impurity correlations, as the former is necessary to induce large attractive interactions between the impurities and thus create the tightly bound states shown in Fig. 4(c). To confirm the crucial role of correlations in the formation of the polaron state we compare with a Bose-Fermi mixture which is separable and therefore uncorrelated. In Fig. 7(a,c) we show the bosonic and fermionic densities with interspecies correlations, and the emergence of the polaron state for increasing $U_{BF}$. We find that for strong coupling $U_{BF}=40$ the density converges to the $U_{\rm BF}\to\infty$ limit, the latter which we obtain by removing the double Bose–Fermi occupancy in the Hilbert space for the DMRG calculations (see Appendix A).
In comparison, we show in Fig. 7(b,d) the densities for the mixture without interspecies correlations, by using a separable state ansatz where we employ a mixed DMRG-exact diagonalization approach for the ground state search (see Appendix A). Here we observe the direct effect of the density-density coupling which confines the fermions in a deeper effective potential which is formed by the bosonic density minima in the centre of the system ($U_{\rm BF}=2$). For large interspecies interactions ($U_{\rm BF}=40$) the system reaches phase separation with the fermions completely localized on two adjacent lattice sites at the center of the trap surrounded by the bosonic density which is entirely depleted at the position of the fermions. At such large coupling, the bosonic translational symmetry is spontaneously broken with an $L$-fold degenerate ground-state labeled by the position of the bipolaron, as suggested by previous works on mixtures [87]. It is important to note that without impurity bath-correlations there are no induced interactions between the fermions ($I_{\rm FF}=I_{\rm{FF},0}$), with their apparent attraction and self-localization being wholly due to the local bosonic effective potential.
As the phase-separated regime in Fig. 7(b,d) can only be reached by neglecting interspecies correlations, it highlights the important role that strong impurity-bath entanglement plays in polaron formation and impurity self-localization. This is typified by the saturation of the von Neumann entropy in Fig. 6(a) which signifies how correlations are locked between the impurities and the bath in the strong coupling regime, and therefore halts further localization of the fermions as shown in Fig. 1(d). This is an example of impurity self-localization suppression enhanced by impurity-bath correlations, an effect which was proposed in Ref. [88] using a variational ansatz in which the impurity-bath correlations are described nonperturbatively. Our work provides further evidence for this effect in discrete lattice systems, and here it is also responsible for the system to completely avoid the phase-separated state.
Figure 7: Comparison of density profiles between a general entangled mixture ansatz (a,c) and a separable (product) ansatz (b,d). (a,b) Bosonic $\expectationvalue{\hat{n}^{\rm B}_{j}}$ and (c,d) fermionic $\expectationvalue{\hat{n}^{\rm F}_{j}}$ density profiles for fixed $U_{\rm BB}=2$, multiple $U_{\rm BF}$ values interactions, including the $U_{\rm BF}\to\infty$ limit for the entangled mixture, and with Periodic(B)-Open(F) BCs.
VI Role of population imbalance
Next, we investigate systems with larger numbers of fermions, $N_{\rm F}>2$, while keeping the number of bosons fixed at $N_{\rm B}=L=30$. In Fig. 8(a-c) we show the density profiles for fixed interactions $U_{\rm BB}=2$ and $U_{\rm BF}\to\infty$, for impurity sizes $N_{\rm F}=6$, $10$ and $14$.
Figure 8: (a,b,c) One-body density profiles for fermions (blue circles) and bosons (red diamonds) with different numbers of impurities at $U_{\rm BB}=2$ and $U_{\rm BF}\to\infty$. (d,e,f) Two-body fermionic density profiles at the same interaction parameters and corresponding impurity numbers.
We see in Fig. 8(a,b), for $N_{\rm F}=6$ and $N_{\rm F}=10$ respectively, that the ground states are $N$-polaron states with density profiles of similar characteristic shape as that in Fig. 2(b). We also note that the fermions become more localized in the centre of the trap as the number of impurities is increased, with the bosonic density consequently decreasing in the same region. For larger numbers of impurities, $N_{\rm F}=14$ in Fig. 8(c), there is no signature of the $N$-polaron state and instead the system is phase separated with the bosonic component is completely depleted at the centre of the trap while the fermions are at unit filling. This shows that the formation of the $N$-polaron state is only present when there is a high interspecies imbalance.
In general, adding impurities to the system at fixed $U_{\rm BB}$ effectively compresses the bath, as the bosons cannot occupy the same place as the fermionic cloud, which is explicitly shown in the bosonic profiles through Fig. 8(a-c). With a denser background, bath-mediated interactions are suppressed which removes the possibility of delocalization of the fermionic cluster and precipitates the collapse of the polaronic state. We can see this in the corresponding two-body density profiles in Fig. 8(d-f), which for $N_{\rm F}=6$ and $N_{\rm F}=10$ (Fig. 8(d,e)) shows qualitatively similar features as in Fig. 4(c). However, it is apparent that the fermions are progressively localized along the center of mass direction ($i=j$) and delocalized along the relative coordinate direction ($i=-j$) as the number of impurities increases and the mediated interactions are diminished. In the phase separation regime at $N_{\rm F}=14$ ((Fig. 8(f)) the two-body density is uniform with equal probability to find a fermion anywhere in the central $14$ lattice sites, with all signature of the polaron state lost.
To study the correlations around the transition to the phase-separated state, the density overlap $\Lambda$ and the mutual information $I_{\rm FF}$ represent the ideal tools in our framework, as the former is zero at phase separation and the latter signals the formation of the $N$-polaron state through the presence of bath-induced correlations when $I_{\rm FF}>I_{\rm FF,0}$, with $I_{\rm FF,0}$ being the mutual information of the $N$ fermions decoupled from the bosonic background ($U_{\rm BF}=0$). We do not use $S_{\rm BF}$ since its calculation is, practically speaking, only possible for very few impurities since the dimension of $\hat{\rho}_{\rm F}$ increases as $L^{N_{\rm F}}\times L^{N_{\rm F}}$. To investigate the miscible-immisible transition for a large number of impurities, $N_{\rm F}=15$, we show in Fig. 9(a) how $I_{\rm FF}/I_{\rm{FF},0}$ changes as a function of interspecies coupling $U_{\rm BF}/U_{\rm BB}$, for fixed boson-boson interactions $U_{\rm BB}=2$, $4$ and $6$.
Figure 9: (a) $I_{\rm FF}$ vs $U_{\rm BF}/U_{\rm BB}$ for $N_{\rm F}=15$ and (b) $I_{\rm FF}$ vs $N_{\rm F}$ for $U_{\rm BF}\to\infty$, with $U_{\rm BB}=2$ (blue dot-dashed), $4$ (purple dashed) and $6$ (orange solid), and Periodic(B)-Open(F) BCs. (c) $\Lambda$ vs $U_{\rm BF}/U_{\rm BB}$ for $N_{\rm F}=15$ and (d) $\Lambda$ vs $N_{\rm F}$ with the corresponding interaction parameters and BCs. We normalize $I_{\rm FF}$ and $\Lambda$ with their non-interacting value $I_{\rm{FF},0}=I_{\rm FF}(U_{\rm BF}=0)$, $\Lambda_{0}=\Lambda(U_{\rm BF}=0)$ and draw gray dotted lines to indicate $U_{\rm BF}=U_{\rm BB}$ (vertical), $I_{\rm FF}=I_{\rm{FF},0}$ and $\Lambda=\Lambda_{0}$ (horizontal) as a reference.
In all cases, we observe that the transition to phase separation is similar to the Open(B)-Open(F) BCs case (Fig. 5(b)), however, while the transition still occurs at larger values of the boson-fermion coupling $U_{\rm BF}>U_{\rm BB}$, the transition point tends to $U_{\rm BF}\rightarrow U_{\rm BB}$ the deeper the bosons are in the MI regime. This is supported by the overlap calculations shown in Fig. 9(c) where $\Lambda/\Lambda_{0}$ decreases from $1$ (miscible) to 00 (immiscible) around the transition point for each $U_{\rm BF}$.
Finally, we investigate the crossover between the $N$-polaron state and the phase-separated state as a function of the number of impurities $N_{\rm F}$. In Fig. 9(b) we show $I_{\rm FF}/I_{\rm{FF},0}$ for different fixed $U_{\rm BB}$ values and in the $U_{\rm BF}\to\infty$ limit whereby we have removed double interspecies occupancies. This figure shows how a larger number of impurities inhibits the bath-mediated interactions which reduce $I_{\rm FF}$ until a critical impurity number $N_{\rm F}^{\rm c}$ is reached at $I_{\rm FF}=I_{\rm{FF},0}$, indicating that the ground state is now phase-separated. This can also be observed in Fig. 9(d) as the overlap goes to zero at the same critical impurity number. We note the important role of the bosonic interactions $U_{\rm BB}$ on the critical impurity number. As can be seen from Fig. 9(b,d), for larger boson-boson interactions, phase separation occurs even for a small number of impurities as boson-boson correlations prevent overlap with the fermionic cloud, analogous to the behavior of $S_{\rm BF}$ at large $U_{\rm BF}$ in Fig. 5(a), therefore hastening the formation of the phase separated state.
VII System size scaling
So far we have shown results for a fixed number of lattice sites ($L=30$), here we explain how these extend to different system sizes while fixing the unit bosonic filling $N_{\rm B}=L$. We focus on the case $U_{\rm BB}=2$ in the limit $U_{\rm BF}\to\infty$ which will signal if the polaron state persists or not for arbitrary large interspecies interactions.
Figure 10: Comparison of fermionic density profiles for $N_{\rm F}=2$ at different system sizes $L=N_{\rm B}$ for $U_{\rm BB}=2$ and in the $U_{\rm BF}\to\infty$ limit. Inset: Critical fermionic density $\rho_{\rm F}^{\rm c}$ for the $N$-polaron to phase separation transition as a function of $1/L$ and a linear extrapolation to the thermodynamic limit ($L\to\infty$).
In Fig. 10 we show the fermionic density profile for $N_{\rm F}=2$ and different system sizes $L$. In all cases, we observe the characteristic density of the delocalized bipolaron state whose width scales with the increasing box size $L$ and shows analogous features as discussed in Sec. III-V.
This raises the question how the polaron state will persist for increasing impurity density and system size, i.e. when does the polaron state transition to the phase separated state as shown in Fig. 8. To answer this we take a similar approach as in Fig. 9(b) where we track the critical density of impurities $\rho_{\rm F}^{\rm c}=N_{F}^{\rm c}/L$ after which $I_{\rm FF}=I_{\rm{FF},0}$, which indicates the phase separation transition for increasing system size $N_{\rm B}=L$. We find this to be a better criterion than the vanishing overlap since there are small contributions to $\Lambda$ in the phase separation regime for finite system sizes which complicate the characterization of the transition. In the inset of Fig. 10 we show the critical fermionic density $\rho_{\rm F}^{\rm c}$ as a function of the inverse system size $1/L$ to study the behavior in the thermodynamic limit ($L\to\infty$). The data shows a linear tendency which converges to a critical density $\rho_{\rm F}^{\rm c}(L\to\infty)=0.3$, hence we conclude that for $U_{\rm BB}=2$ and $\rho_{\rm F}<\rho_{\rm F}^{\rm c}(L\to\infty)$ the $N$-polaron state can emerge for arbitrary system sizes.
VIII Conclusions
@@
We have studied the formation of a $N$-polaron state, where fermionic impurities are tightly bound together by a unit-filled bosonic gas in a one-dimensional optical lattice. We have shown that a superfluid-insulator transition is crucial for polaron formation, as when the Mott phase is broken and the background becomes a superfluid, correlations with the bosonic bath bind fermions together through strong mediated attractive interactions. This allows for the redistribution of the polaron across the superfluid bath and prevents phase separation, similar to Ref. [88] and recently observed in DMRG simulations [89].
This also suggests that the presence of strong impurity-bath correlations could inhibit the formation of the self-pinned state [90], which relies on a mean-field interspecies coupling for the particles to self-localize and reorganize into a periodic structure. While the separable ansatz used at the end of Sec. V could realize an analog of the self-pinned state in our lattice model, using DMRG for the full Hamiltonian resulted in different physics since the bath-mediated correlations favor the delocalized polaronic state. However, in our results, we do find promising evidence that the self-pinned state begins to emerge when the impurities self-localize before the MI-SF transition, as explained in Sec. IV. We then expect that the self-pinned state could emerge by damping mediated correlations in the bath which can be achieved by exploring regimes of larger bosonic density, similar to the phase separation observed for large population imbalance discussed in Sec. VI, or by tuning the mass ratio in the mixture.
We have also shown that population imbalance is key for $N$-polaron formation, and that suitable tuning of bath intraspecies interactions and correlations could allow the creation of large and stable strongly-entangled $N$-polaron states. Our analysis allowed us to determine a fermionic density threshold below which the polaron emerges that could be of use in possible experimental realizations. Indeed, a recent similar work has explored how the filling fraction and boson-boson interactions can affect bipolaron formation elucidating complementary results to our own [91]. Future studies could also investigate the role of particle statistics inside the polaronic state, analog to previous work with bosonic impurities using exact diagonalization [49, 52]. Moreover, the use of mixed BCs on other well-known impurity systems could lead to novel phenomena by blocking or enhancing boundary effects.
@@ Acknowledgements.
Appendix A Ground-state search using DMRG
To find the ground state of the Hamiltonian (1)-(4), the Density Matrix Renormalization Group (DMRG) technique [92] is used in the tensor network formalism [76] with the help of the TeNPy library [93]. As it is necessary to truncate the local bosonic Hilbert space, the soft-core approximation is used for the maximum number of bosons per site $N_{\rm B}^{\rm{max}}=5$, which is large enough to capture the essential physics for $U_{\rm BB}\geq 2$ [94, 81, 95, 96] and results in a local basis of size $12$ with all of the possible combinations of fermionic and bosonic occupations under the assumed approximation. In the $U_{\rm BF}\to\infty$ approximation the local basis is reduced to $7$ elements by removing the states with a boson and a fermion on the same site.
Given that a periodic boundary condition adds a long-range hopping between the edges of the one-dimensional lattice chain, for the construction of the matrix product state (MPS) we reorder the system as $\{s_{1},s_{L},s_{2},s_{L-1},...\}$, where $s_{j}$ represents the $j$-th site in the original ordering. This increases the nearest-neighbor hopping to the next-to-nearest neighbor but reduces the long-range hopping at the edges to a nearest-neighbor term. The advantage of this transformation is that it preserves locality in the open one-dimensional system represented by the MPS which upholds the validity of the area law scaling of a subsystem’s entanglement entropy with the size of its boundary, crucial for the efficiency of MPS algorithms [97]. On the other hand, alternatives such as a pure periodic MPS [76] where an additional bond closes the MPS by its edges require a more elaborate framework [98, 99, 100] since it is not possible to split the system in two with a singular value decomposition, a necessary operation for most of the MPS algorithms, hence we did not consider this alternative method to be able to use the available software for standard MPS [93].
During the DMRG algorithm, a maximum truncation error of the many-body Hilbert space of $10^{-14}$ is set; based on this criteria, the library optimizes the bond dimension $\chi$ along the chain with a fixed upper bound $\chi_{\rm{max}}$. To improve the algorithm’s convergence, we ramp up $\chi_{\rm{max}}$ by 100 for every 10 sweeps until it reaches a maximum cap of 3000. The program optimizes the wave function for each simulation until the ground state energy has an error less than $10^{-7}$ and the entanglement entropy at the middle of the chain less than $10^{-3}$.
We also consider the case of a separable wave function between species. To tackle this problem, we split the bosonic (2) and fermionic (3) Hamiltonians and minimize the energy iteratively, one component at a time, while taking into account the density-density coupling (4). We use an MPS ansatz for the bosonic species to capture the correlations of the Bose Hubbard model and apply a DMRG ground-state search; for each sweep, we fix the impurity density, while between sweeps, the bosonic background is fixed, and the fermionic Hamiltonian is solved exactly, given that it is quadratic. Since the local basis for the Bose Hubbard model is smaller than the mixed case ($N_{\rm B}^{\rm{max}}+1=6$ states), we only require bond dimensions of $\chi_{\rm max}\approx 300$. The rest of the parameters are kept the same as in the complete model.
Appendix B Bosonic correlations in the SF-MI transition
Self-localization in this system emerges because of the difference in the spread of correlations between MI and SF backgrounds. This appendix intends to characterize this for the Periodic(B)-Open(F) BCs case as the focus of this work, by using the bosonic correlation function $\widetilde{\rho}^{(1)}_{\rm B}$ where we average the corresponding reduced density matrix at a given distance:
$\widetilde{\rho}^{(1)}_{\rm B}(d)=\frac{\sum_{j,k}^{L}\delta_{\min\left(|j-k|,L-|j-k|\right),d}\expectationvalue{\hat{b}^{\dagger}_{k}\hat{b}_{j}}}{\sum_{j,k}^{L}\delta_{\min\left(|j-k|,L-|j-k|\right),d}}.$ (11)
Here, the Kronecker delta signals the pair of sites at a distance $d$ from each other, taking into account the periodic boundary conditions of the bosonic system.
We plot $\widetilde{\rho}^{(1)}_{\rm F}(d)$ in Fig. 11 for $U_{\rm BB}=6$ and different $U_{\rm BF}$ values. The maximum of the relative distance between the fermions signals the MI-SF transition ($U_{\rm BF}\approx 4$, see Fig. 1(f)), so we consider representative values of interspecies interaction below, $U_{\rm BF}=0$ and $U_{\rm BF}=3$, and above, $U_{\rm BF}=6$ and $U_{\rm BF}=9$, this transition. For $U_{\rm BF}=0$ and $U_{\rm BF}=3$ the decay of the correlation function is exponential, except for finite size effects at large $d$, a behavior characteristic of an insulating state which in turn suppresses correlation between the fermions.
Figure 11: Averaged correlation function $\widetilde{\rho}^{(1)}_{\rm B}$ vs distance $d$ for $U_{\rm BB}=6$ and $U_{\rm BF}=0$, $3$, $6$ and $9$ with Periodic(B)-Open(F) BCs.
Then, for interactions $U_{\rm BF}=6$ and $U_{\rm BF}=9$ larger than the transition critical value, long-range correlations in the form of an algebraic decay indicate the SF background transition where boson-mediated fermionic entanglement can be generated.
Appendix C Quasiparticle residue
A key concept in the analysis of polaron physics is the quasiparticle weight [17], defined as the squared overlap between the coupled ($U_{\rm BF}=0$) and uncoupled ($U_{\rm BF}\neq 0$) states. We define the quasiparticle weight through the overlap of the full wavefunctions
$Z=|\innerproduct{U_{\rm BF}=0,U_{\rm BB}}{U_{\rm BF},U_{\rm BB}}|^{2}\,,$ (12)
with the interacting wave function |$\ket{U_{\rm BF},U_{\rm BB}}$ and the initial decoupled state |$\ket{U_{\rm BF}=0,U_{\rm BB}}$. In Fig. 12, we show this quantity as a function of the interspecies interaction for multiple $U_{\rm BB}$ values, all of which show the convergence of $Z\to 0$ as we increase $U_{\rm BF}$. We would like to stress that this quantity is strongly affected by the larger bosonic component, which dominates the overlap as $N_{\rm B}\gg N_{\rm F}$. In the regime of strong impurity-bath interaction $U_{\rm BF}/U_{\rm BB}\gg 1$, the impurities strongly perturb the bosonic bath and drive it to orthogonality. This is also accelerated around a quantum critical point in the system, which we have when the impurities induce the MI-SF transition and the fermions collapse into the bound polaronic state. Crucially, the overlap will decay with increasing number of bosons, $Z\sim N_{\rm B}^{-\alpha}$, as described by the orthogonality catastrophe [101]. The vanishing of the quasiparticle residue has also been explored previously for Bose-polarons [102] and is a characteristic of 1D systems, so we see that the quasi-particle residue vanishing in our case is to be expected.
Figure 12: Quasiparticle weight $Z$ as a function of the interspecies interaction $U_{\rm BF}$ scaled by a fixed $U_{\rm BB}=2$, $4$ and $6$. $Z=0$ is shown as a dotted line for reference.